# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ 4 Clinical Applications
As shown in Figure 5, this section discusses the clinical applications of LLMs. Each subsection contains a specific application and discusses how LLMs perform this task.
@@ 4.1 Medical Diagnosis
Medical diagnosis involves the medical practitioner using objective medical data from tests and self-described subjective symptoms to conclude the most likely health problem occurring in the patient AImed . It is important to diagnose a patient in an accurate and timely manner because the effectiveness of treatment for most diseases is time-sensitive. A missed or inaccurate diagnosis often has negative consequences, ranging from a minor inconvenience to death. For example, breast cancer has the highest mortality rate in the world because many communities lack trained personnel to perform proper checks gao2023designing . It may increase the accessibility of professional healthcare AImed ; thirunavukarasu2023large to incorporate LLMs into the medical diagnosis pipeline.
@@ Discussion
One distinct limitation of using LLMs as the sole tool for medical diagnosis is the heavy reliance on subjective text inputs from the patient. Since LLMs are text-based, they lack the inherent capability to analyze medical diagnostic imagery. Given that objective medical diagnoses frequently depend on visual images, LLMs are often unable to directly conduct diagnostic assessments as they lack concrete visual evidence to support disease diagnosis liu2021PPKED . However, they can help with diagnosis as a logical reasoning tool for improving accuracy in other vision-based models. One such example is ChatCAD wang2023chatcad , where images are first fed into an existing computer-aided diagnosis (CAD) model to obtain tensor outputs. These outputs are translated into natural language, which is subsequently fed into ChatCAD to summarize results and formulate diagnoses. ChatCAD achieves a recall score of 0.781, significantly higher than that (0.382) of the state-of-the-art task-specific model. We note that although the more recent GPT-4V(ision) is capable of interpreting images in the general domain, its extension to the medical image domain is yet to materialize. Nevertheless, all aforementioned methods of implementing LLMs rely on image transformation into text beforehand.
@@ 4.2 Formatting and ICD-Coding
The international classification of diseases (ICD) world1978international is a method of standardizing diagnostic and procedural information of a clinical session. These ICD codes are recorded in the patient’s EHRs every doctor visit. They are also used for tracking health metrics, treatment outcomes, and billing. There is a need to automate the ICD labeling process because its manual entry process is very time consuming for doctors.
LLMs can help automate ICD coding by extracting medical terms from clinical notes and assigning corresponding ICD codes ong2023applying ; dong2022automated . For example, PLM-ICD huang2022plm is a multi-class classification model fine-tuned for automatic ICD coding. The base model used in PLM-ICD is domain-specific with medicine-specific knowledge to enhance the ability to understand medical terms. PLM-ICD uses segment pooling, the algorithm that divides long input texts into shorter representations using LLMs when the input surpasses the maximum allowable length. Lastly, it relates the encoding to the augmented labels to output ICD codes for each clinical input. PLM-ICD produced a higher AUC score than previous state-of-the-art lightweight models huang2022plm .
@@ Discussion
One important issue with LLMs is the potential biases and hallucinations. In particular, traditional multi-label classification models can easily constrain their outputs to a predefined list of (usually >1000) ICD candidate codes through a classification neural network. In contrast, generative LLMs struggle to effectively handle this extensive list through prompting techniques. As a result, the LLM may assign an ICD code that is not in the candidate list or a non-existent ICD code to the input text. It leads to confusion when interpreting medical records yang2022large and is therefore crucial to establish a proactive mechanism to detect and rectify errors before they enter patient EHRs.
@@ 4.3 Clinical Report Generation
Clinical reports, such as radiology reports wu2023towards , discharge summaries patel2023chatgpt , discharge instruction liu2022retrieve , and patient clinic letters ali2023using , refer to standardized documentation that healthcare workers complete after each patient visit liu2019reporting . A large portion of the report is often medical diagnostic results. It is typically tedious for overworked clinicians to write clinical reports, and thus they are often incomplete or error-prone. Adopting LLMs for clinical report generation can provide an objective means of increasing their completeness and accuracy, as well as reducing clinical workload. The generated report is intended to be a document that the clinician can review, modify, and approve as necessary patel2023chatgpt ; ali2023using ; liu2022retrieve .
Meanwhile, LLMs can be used intuitively as a summarization tool wang2023chatcad to help with clinical report generation. Given the diagnosis as input, an LLM can use its capabilities of text summarization to produce a clear and concise conclusion. In this instance, LLMs do not directly contribute to improving the accuracy of the conclusion. Instead, they act as a tool for the work that otherwise would have been undertaken by doctors.
Another popular utilization of LLMs for generating clinical reports relies on a vision-based model or manual input acting as a precursor in the pipeline wang2023chatcad ; wu2023towards ; wu2023can . This precursor step analyzes the input medical image and feeds the LLM with some type of annotation. The LLM uses the annotation alongside some other text prompt (e.g. report format, additional input by medical personnel) to generate an accurate and fluent report that follows the parameters. This has been shown to greatly reduce the workload on doctors wang2023chatcad .
@@ Discussion
Although LLMs have generated more complete and accurate clinical reports than human counterparts van2023clinical , they are still susceptible to hallucinations as well as tend to approach inputs with a literal view instead of an assumption-based perspective that is often taken by human doctors. Additionally, human-written reports are generally more concise than reports generated by LLMs van2023clinical .
@@ 4.4 Medical Robotics
Medical robotics are revolutionizing healthcare, offering precision in tasks ranging from medication distributioncheng2023multi to diagnostic analyses and multi-agent surgical planning xia2023nested . These intelligent systems enhance the medical workforce, bridge gaps in staffing, and perform tasks that extend human physical limits.
@@ Discussion
Some challenges with implementing medical robotics are similar to those with collaborative robots (cobots), as both cases involve robots operating alongside humans that require trust between human and computeremaminejad2023trust . Different from cobots, when implementing LLM into algorithms for route planning and robotic motion, errors in judgment due to the effects of bias and hallucination may occur. This concern can be mitigated by the better ability of traditional robots to inhibit damage than cobots. The low prediction accuracy score (< 0.5) of current LLM-powered medical robots means they are still unsuitable for implementation in real clinical settings cheng2023multi ; xia2023nested ; robinson2023brief , which also indicates research opportunities.
@@ 4.5 Medical Language Translation
There are two main areas of medical language translation; the translation of medical terminology from one language to another noll2023machine and the translation of professional medical dialogue into expressions that are easily interpreted by non-professional personnel karabacak2023advent . Both areas are important for seamless communication between different groups. The current language barrier to global collaboration in both research and medical techniques can be largely reduced with the help of LLMs. Machine translation has been proven to be 7% more accurate than traditional human services karabacak2023advent . Language translation also improves accuracy in education resources and research articles, making knowledge accessible worldwide ahn2023impending . For the second use case, LLMs can improve medical education by identifying the skill levels of students and providing corresponding terminology and structures they will understand. This application will also help patients understand professional medical speech karabacak2023advent .
@@ Discussion
One ethical consideration of using LLMs to perform medical translation is the potential for discriminatory verbiage to be inserted inadvertently into the output. Such verbiage is difficult to prevent due to the nature of the pipeline. This may cause miscommunications and even have legal consequences. Additionally, potential misinformation caused by translation errors may confuse patients and, in the worst case, may cause them to unintentionally inflict harm to themselves karabacak2023advent .
@@ 4.6 Medical Education
LLMs can be incorporated into the medical education system in different ways, including facilitating study through explanations, aiding in language translation, answering questions, assisting with medical exam preparation, and providing Socratic-style tutoring qiu2023large ; biri2023assessing . It has been suggested that medical education can be augmented by generating scenarios, problems, and corresponding answers by an LLM. Students will gain a richer educational experience through personalized study modules and case-based assessments, encountering a wider array of challenges and scenarios beyond those found in standard textbooks karabacak2023advent . LLMs can also generate feedback on the students’ responses to practical problems, allowing students to know their areas of weakness in real-time. Inherently, these will better prepare these medical students for the real world since they would have been exposed to more scenariosahn2023impending .
@@ Discussion
Potential downsides of using LLMs in medical education include the current lack of ethical training and biases in training datasetsabd2023large . In addition, misinformation (e.g. hallucination) presents a challenge in utilizing LLMs for medical education.
@@ 4.7 Mental Health Support
Mental health support involves both diagnosis and treatment. For example, depression is treated through a variety of psychotherapies, including cognitive behavior therapy, interpersonal psychotherapy, psychodynamic therapy, etc. vaidyam2019chatbots . Many of these techniques are primarily dominated by patient-doctor conversations, with lengthy treatment plans that are cost-prohibitive for many. The ability of LLMs to serve as conversation partners and companions may significantly lower the barrier to entry for patients with financial or physical constraints stock2023tell , increasing the accessibility to mental health treatments liu2023chatcounselor . There have been various research works on the effects of incorporating LLMs into the treatment plan liu2023chatcounselor ; de2023benefits .

### v5

@@ 4 Clinical Applications
As shown in Figure 5, this section discusses the clinical applications of LLMs. Each subsection contains a specific application and discusses how LLMs perform this task. Table 3 summarizes the guidelines on how to select, build, and evaluate medical LLMs for various clinical applications.
@@ 4.1 Medical Diagnosis
Medical diagnosis involves the medical practitioner using objective medical data from tests and self-described subjective symptoms to conclude the most likely health problem occurring in the patient AImed. This heavily relies on the synthesis and interpretation of vast amounts of information from various sources, including patient medical histories, clinical data, and the latest medical literature. The advent of large language models has opened up new opportunities for enhancing medical diagnostic processes. These advanced natural language processing models can rapidly process and comprehend massive volumes of medical data, literature, and legal guidelines, potentially aiding healthcare professionals in making more informed and legally sound diagnostic decisionsAImed; thirunavukarasu2023large.
Guideline
Dr. Knowsgao2023leveraging extracts knowledge graph from the unified medical language system (UMLS) bodenreider2004unified for identifying likely diagnoses from patient medical records. It is then further fine-tuned T5 models using these diagnoses as prompts and leveraged zero-shot prompting for ChatGPT, demonstrating improved diagnosis prediction accuracy by utilizing the knowledge graph. An alternative is DDx PaLM-2mcduff2023towards, which builds on Google’s PaLM-2 and is further fine-tuned with medical datasets for zero-shot prompting. It can engage in conversations and collaborate with clinicians, allowing them to interactively identify potential diagnoses for complex medical conditions while improving the doctors’ diagnostic reasoning abilities through the interactive process. Recent advancement in multimodal LLMs allows AI to understand more complex medical knowledge and data with multimodality zhao2023clip. Notable examples include Med-Flamingomoor2023med, LLaVA-Medli2024llava, and Med-Geminisaab2024capabilities. Med-Flamingomoor2023med builds on Google’s vision-language model Flamingo and is pret-rained with medical image-text data. LLaVA-Medli2024llava, proposed by Microsoft, builds on the LLaVA model and is trained with a two-stage medical concept alignment and medical instruction tuning using 660k samples. Google’s powerful Med-Geminisaab2024capabilities builds upon the Gemini foundation and is further fine-tuned with large-scale medical multimodal data, highlighting its capabilities in multimodal understanding and long-context processing. In training and evaluating these models, commonly utilized datasets include the MIMIC series, such as MIMIC-IIIjohnson2016mimic and MIMIC-IVjohnson2020mimiciv discharge summaries, as well as question answering datasets such as MultiMedQAsinghal2023large, which constructed from various resources covering patient cases from PubMed articles and textbooks. For multimodal data, VQA-RAD (radiology)lau2018dataset, SLAKE (radiology)liu2021slake, and PathVQA (pathology)he2020pathvqa are frequently employed. Most benchmarking efforts involve both quantitative evaluation metrics and human evaluations. These models have demonstrated their effectiveness and the potential for substantial improvements in medical diagnosis tasks.
@@ Discussion
One distinct limitation of using LLMs as the sole tool for medical diagnosis is the heavy reliance on subjective text inputs from the patient. Since LLMs are text-based, they lack the inherent capability to analyze medical diagnostic imagery. Given that objective medical diagnoses frequently depend on visual images, LLMs are often unable to directly conduct diagnostic assessments as they lack concrete visual evidence to support disease diagnosis liu2021PPKED. However, they can help with diagnosis as a logical reasoning tool for improving accuracy in other vision-based models. One such example is ChatCAD wang2023chatcad, where images are first fed into an existing computer-aided diagnosis (CAD) model to obtain tensor outputs. These outputs are translated into natural language, which is subsequently fed into ChatCAD to summarize results and formulate diagnoses. ChatCAD achieves a recall score of 0.781, substantially higher than that (0.382) of the state-of-the-art task-specific model. Nevertheless, all the aforementioned methods of implementing LLMs cannot directly process images; instead, they either rely on transforming images into text beforehand or rely on an external separate vision encoder to embed images.
@@ 4.2 Formatting and ICD-Coding
The international classification of diseases (ICD) world1978international is a method of standardizing diagnostic and procedural information of a clinical session. These ICD codes are recorded in the patient’s EHRs every doctor visit. They are also used for tracking health metrics, treatment outcomes, and billing. There is a need to automate the ICD labeling process because its manual entry process is very time-consuming for doctors. Formatting and ICD-Coding usually involve entity extraction, relation extraction, text generation, and information retrieval. LLMs can help automate ICD coding by extracting medical terms from clinical notes and assigning corresponding ICD codes ong2023applying; dong2022automated.
Guideline
For example, PLM-ICDhuang2022plm builds upon the RoBERTa model liu2019roberta, fine-tuning it specifically for ICD coding and achieving strong performance on 70,539 notes from the MIMIC-II and MIMIC-III datasets johnson2016mimic, as evaluated by accuracy. The base model used in PLM-ICD is domain-specific with medicine-specific knowledge to enhance the ability to understand medical terms. PLM-ICD uses segment pooling, the algorithm that divides long input texts into shorter representations using LLMs when the input surpasses the maximum allowable length. Lastly, it relates the encoding to the augmented labels to output ICD codes for each clinical input. PLM-ICD produced a higher AUC score than previous state-of-the-art lightweight models huang2022plm. DRG-LLaMAwang2024drg leverages the LLaMA model and applies parameter-efficient fine-tuning techniques, such as LoRA, to adapt the model to this task. ChatICDliu2023chaticd and LLM-codexyang2023surpassing both utilize the ChatGPT model with prompts for ICD coding. However, LLM-codexyang2023surpassing takes this a step further by training an LSTM model on top of the ChatGPT responses, demonstrating its strong performance. ICD coding can be formulated as a multi-label classification task, and most work in this area utilizes the MIMIC-III dataset for training and evaluation. Models are typically assessed based on their F1 score, AUC, and Precision@k, considering either the top 50 most frequent labels or the full label set.
@@ Discussion
One challenge while deploying LLMs for clinical coding is the potential biases and hallucinations. In particular, traditional multi-label classification models can easily constrain their outputs to a predefined list of (usually >1000) ICD candidate codes through a classification neural network. In contrast, generative LLMs can suffer from major hallucinations while the input text is lengthy. As a result, the LLM may assign an ICD code that is not in the candidate list or a non-existent ICD code to the input text. It leads to confusion when interpreting medical records yang2022large and is, therefore, crucial to establish a proactive mechanism to detect and rectify errors before they enter patient EHRs.
@@ 4.3 Clinical Report Generation
Clinical reports, such as radiology reports, discharge summaries, and patient clinic letters, refer to standardized documentation that healthcare workers complete after each patient visit liu2019reporting. Therefore, clinical report generation usually involves text generation/summarization, and information retrieval. A large portion of the report is often medical diagnostic results. It is typically tedious for overworked clinicians to write clinical reports, and thus they are often incomplete or error-prone. Meanwhile, LLMs can be used intuitively as a summarization tool to help with clinical report generation. In this instance, LLMs act as an assistant tool for clinicians which helps improve efficiency and reduce potential errors in lengthy reports patel2023chatgpt; ali2023using; liu2022retrieve.
Another popular approach to generating clinical reports using LLMs involves incorporating a vision-based model to provide complementary information wang2023chatcad; wu2023towards; wu2023can. The vision model analyzes the input medical image and generates an annotation, which serves as a direct and supplementary input to the LLM alongside additional text prompts. By leveraging the combination of visual and textual information, the LLM produces accurate and fluent reports that adhere to the specified parameters and structure.
Guideline
For radiology report generation, general medical vision-language models like Med-Geminisaab2024capabilities, LlaVA-Medli2024llava, and Med-Flamingomoor2023med can be utilized, but models trained specifically on radiographs, such as ChatCADwang2023chatcad, MAIRA-1hyland2023maira and RadFMwu2023towards, have shown superior performance than the general medical ones. Alternatively, language models like ImpressionGPTma2024iterative leverage textual data for report summarization, employing dynamic prompt generation and iterative optimization. RadAdaptvan2023radadapt systematically evaluates various language models and lightweight adaptation methods, achieving optimal performance through pre-training on clinical text and parameter-efficient fine-tuning with LoRA, investigating the impact of few-shot prompting. In terms of evaluation, most work uses MIMIC-III or MIMIC-IV notes for training and evaluation, as it is the largest publicly available free-text EHRs. Common automatic evaluation metrics includes lexical methods such as BLEUpapineni2002bleu, ROUGE lin2004rouge, METEORbanerjee2005meteor, semantic-based methods such as BERTScorezhang2019bertscore, and radiology-specific metrics such as CheXbert similaritysmit2020chexbert, RadGraphjain2021radgraph, RadCliQyu2023evaluating.
@@ Discussion
While LLMs have demonstrated the ability to generate clinical reports that are more comprehensive and precise than those written by human counterpartsvan2023clinical, they still face challenges in terms of hallucinations and literal interpretation of inputs, lacking the assumption-based perspective often employed by human doctors. Moreover, LLM-generated reports tend to be less concise compared to human-written ones. The evaluation of LLMs in this domain is particularly challenging due to the specialized nature of the content and the generative nature of the task. Current automatic evaluation methods for clinical report generation primarily focus on lexical metrics, which can lead to biased and inaccurate assessments of the contextual information present in the reports xie2023faithful. For instance, consider two sentences with similar meanings but different wordings: “The patient’s blood glucose level is within normal limits” and “The patient does not exhibit signs of hyperglycemia”. While both convey the absence of hyperglycemia, lexical evaluation metrics may struggle to accurately capture their semantic equivalence, as they rely on direct word-level comparisons. This discrepancy highlights the need for more sophisticated evaluation techniques that can account for the nuances and variations in expressing clinical information. Developing evaluation methods that go beyond surface-level similarities and consider the underlying medical context is crucial for ensuring the reliability and usefulness of LLMs in generating clinical reports.
@@ 4.4 Medical Robotics
Medical robotics are revolutionizing healthcare, offering precision in ultrasound scanning, diagnostic analyses, and multi-agent surgical planning xu2024enhancing; killeen2024take; moghani2024sufia. These intelligent systems enhance the medical workforce, bridge gaps in staffing, and perform tasks that extend human physical limits. Incorporating LLMs, robots interpret environmental data to navigate and execute complex tasks.
@@
Guideline
SuFIA moghani2024sufia incorporates the strong reasoning capabilities of LLMs with perception modules to implement high-level planning and low-level control of a robot for surgical sub-task execution, with the current best results obtained via API calls to GPT-4 Turbo. While measures are incorporated to improve safety and reliability, deploying autonomous or semi-autonomous robotic surgical assistants (RSAs) in real-world scenarios still has the potential risks from unexpected circumstances. Another work UltrasoundGPT xu2024enhancing proposes an ultrasound-bodied intelligence system equipping ultrasound robots with LLMs and domain knowledge xu2024enhancing. It designs an ultrasound operation knowledge database for the LLM to enable precise motion planning. A dynamic scanning strategy based on prompt engineering allows LLMs to adjust motion planning during procedures. Their system improves ultrasound scan efficiency and quality from verbal commands, contributing to non-invasive diagnostics and streamlined workflows. Another work killeen2024take attempts to interpret domain-specific language in X-ray-guided surgery. It consists of a minimal protocol enabling an LLM, i.e., GPT-4, to control a robotic X-ray system, namely the Brainlab Loop-X device.
@@ Discussion
Some challenges with implementing medical robotics are similar to those with collaborative robots (cobots), which are designed to work alongside and interact with human workers in shared spacesemaminejad2023trust. Integrating LLMs into medical robotics algorithms for route planning and motion control poses a critical challenge due to the risk of errors and biases inherent in LLMs. The complex and dynamic nature of shared human-robot workspaces may lead to LLM-powered cobots misjudging human intentions or making inappropriate decisions, posing safety risks. In contrast, traditional industrial robots, designed to operate at high speeds in controlled, human-free environments, could better withstand and contain errors or malfunctions due to their rigid and powerful designs, effectively mitigating human injury risks. Future research opportunities could explore safety features for cobots, such as sophisticated sensing technologies and physical design constraints, which aim to minimize the occurrence and consequences of judgment errors related to LLMs in shared human-robot environments weerarathna2023human; moglia2021systematic; xia2023nested.
@@ 4.5 Medical Language Translation
There are two main areas of medical language translation; the translation of medical terminology from one language to another noll2023machine and the translation of medical dialogue for ease of interpretation by non-professional personnel karabacak2023advent. Both areas are important for seamless communication between different groups. The language barrier to global collaboration in both research and medical techniques can be largely reduced with the help of LLMs. Language translation also improves accuracy in education resources and research articles, making knowledge accessible worldwide ahn2023impending.
Guideline
Medical mT5garcia2024medical, Apollowang2024apollo, and BiMediXpieri2024bimedix are multilingual large language models in the medical domain. Medical mT5garcia2024medical, which is based on multilingual T5 (mT5) with 738 million / 3 billion parameters, is trained on 4.5 billion tokens consisting of various languages languages, i.e., English, French, Italian, and Spanish. Apollowang2024apollo supports English, Chinese, French, Spanish, Arabic, and Hindi based on the Qwen model at various relatively small sizes (i.e., 0.5B, 1.8B, 2B, 6B, and 7B), achieving the best performance among models of equivalent size. BiMediXpieri2024bimedix is a bilingual medical mixture-of-experts language model for English and Arabic, proposing a semi-automated English-to-Arabic translation pipeline with human refinement for high-quality translations. For medical translation to lay language, a work aims to enhance the performance of language models in biomedical abstractive summarisation by aggregating knowledge from external papers cited within the source articletang2023improving. It proposes a novel attention-based citation aggregation model that integrates domain-specific knowledge from citation papers, allowing neural networks to generate summaries by leveraging both the paper content and relevant knowledge from citation paperstang2023improving. Another work introduces Retrieval-Augmented Lay Language (RALL) generation with a large and broad-ranging 63k lay language generation pairs from 12 journals, intuitively fitting the need for external knowledge beyond expert-authored source documentsguo2024retrieval. It also evaluates the ability of both an open-source LLaMA-2 and closed-source GPT-4 in background explanation, with and without retrieval augmentation.
@@ Discussion
In both translation and simplification tasks, misinterpretation is a common occurrence that can have damaging consequences. In developing and deploying medical translation and simplification platforms, developers should prioritize professional datasets, such as textbooks and peer-reviewed journals for medical knowledge recall. This way, it will be less likely for misinformation from unreliable web sources to skew the output chen2023can. Another ethical consideration of using LLMs to perform medical translation is the potential for discriminatory verbiage to be inserted inadvertently into the output. Such verbiage is difficult to prevent due to the nature of the pipeline. This may cause miscommunications and even have legal consequences. karabacak2023advent.
@@ 4.6 Medical Education
LLMs can be incorporated into the medical education system in different ways, including facilitating study through explanations, aiding in language translation, answering questions, assisting with medical exam preparation, and providing Socratic-style tutoring qiu2023large; biri2023assessing. Therefore, medical education could involve text generation, text simplification, semantic textual similarity, information retrieval, and etc. It has been suggested that medical education can be augmented by generating scenarios, problems, and corresponding answers by an LLM. Students will gain a richer educational experience through personalized study modules and case-based assessments, encountering a wider array of challenges and scenarios beyond those found in standard textbooks karabacak2023advent. LLMs can also generate feedback on student responses to practical problems, allowing students to know their areas of weakness in real time. Inherently, these will better prepare these medical students for the real world since they would have been exposed to more scenariosahn2023impending.
@@
Guideline
Large language models such as ChatGPTopenai2023chatgpt, and Med-Geminisaab2024capabilities are increasingly demonstrating their potential in medical education applications. These powerful models, trained on vast amounts of medical data, offer capabilities in knowledge synthesis, question answering, and content generation that can augment traditional teaching methods. For instance, ChatGPTopenai2023chatgpt can provide explanations and clarifications on complex medical concepts, facilitating self-study and reinforcing understanding. Med-Geminisaab2024capabilities, a multimodal model, can analyze medical images and generate detailed reports, aiding in the training of diagnostic skills. Institutions are exploring the integration of these language models into curricula, leveraging their strengths while ensuring proper oversight and ethical considerations. As this technology continues to advance, it holds promise for enhancing the efficiency and accessibility of medical education while complementing human expertise.
@@ Discussion
Potential downsides of using LLMs in medical education include the current lack of ethical training and biases in training datasetsabd2023large. These biases, if not addressed, can propagate through the generated outputs, reinforcing stereotypes and potentially leading to discrimination in medical education. The lack of explicit ethical training during LLM development may also result in the generation of content that does not align with the ethical principles and guidelines of the medical profession, such as promoting unethical practices or violating patient privacy.
Furthermore, the risk of misinformation, particularly in the form of hallucinations, presents a challenge in utilizing LLMs for medical education. LLMs can generate plausible-sounding but factually incorrect information, which can mislead students and healthcare professionals if relied upon without proper verification. This can lead to the propagation of misconceptions, inappropriate treatment strategies, or misdiagnosis rawte2023survey. To mitigate these risks, it is essential to establish rigorous fact-checking and validation processes and emphasize the importance of critical thinking, evidence-based practice, and the verification of information from multiple reliable sources in medical education.
@@ 4.7 Mental Health Support
Mental health support involves both diagnosis and treatment. For example, depression is treated through a variety of psychotherapies, including cognitive behavior therapy, interpersonal psychotherapy, psychodynamic therapy, etc. vaidyam2019chatbots. Many of these techniques are primarily dominated by patient-doctor conversations, with lengthy treatment plans that are cost-prohibitive for many. The ability of LLMs to serve as conversation partners and companions may lower the barrier to entry for patients with financial or physical constraints stock2023tell, increasing the accessibility to mental health treatments liu2023chatcounselor. There have been various research works and discussions on the effects of incorporating LLMs into the treatment plan liu2023chatcounselor; de2023benefits; hua2024large.
@@
Guideline
Development and deployment of LLMs targeted at mental health support can start with an existing LLM. Instead of pre-training or fine-tuning on general medical data, it is often better to use medical question and answer data as most of the LLM’s work will be talking to the patient, which involves back-and-forth conversation in the format of question and answering lai2023psy. PsyChat qiu2023psychat is a client-centric LLM dialogue system that provides psychological support comprising five modules: client behavior recognition, counselor strategy selection, input packer, response generator, and response selection. Specifically, the response generator is fine-tuned with ChatGLM-6B with a vast dialogue dataset. Through both automatic and human evaluations, the system has demonstrated its effectiveness and practicality in real-life mental health support scenarios. ChatCounselor is designed to provide mental health support. It initializes from Vicuna and fine-tunes from an 8k size instruct-tuning dataset collected from real-world counseling dialogue examples liu2023chatcounselor. Psy-LLM is an LLM aimed to be an assistive mental health tool to support the workflow of professional counselors, particularly to support those who might be suffering from depression or anxiety lai2023psy. Another work presents a comprehensive evaluation of prompt engineering, few-shot, and fine-tuning techniques on multiple LLMs in the mental health domainxu2024mental. The results reveal that fine-tuning on a variety of datasets can improve LLM’s capability on multiple mental-health-specific tasks across different datasets simultaneouslyxu2024mental. The work also releases their model Mental-Alpaca and Mental-FLAN-T5 as open-source LLMs targeted at multiple mental health prediction tasksxu2024mental.